OpenStudy (anonymous):

2x+y=46 y=x+4 solve equation

OpenStudy (blurbendy):

well we know that y = x+4, so let's put that into the first equation 2x + _________ = 46 fill in the blank with what you know

OpenStudy (anonymous):

2x+y=46 y=x+4 2x+y=46 -2x+2y=8 ---------- 3y=54 y=18 18=x+4 -4 x= 14 (14, 18)

OpenStudy (sk8rgirl12):

It's trial and error. It's not 9. It's not 10 Not 11, or 12, nor 13. I agree with Haiqa. x=14, y=18